OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. whole numbers B. negative integers C. integers D. rational numbers

OpenStudy (matlee):

A and C

OpenStudy (anonymous):

/\ hes right

OpenStudy (anonymous):

is that correct everyone

OpenStudy (matlee):

YES

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no 0 and 1 are whole numbers but 0-1=-1 is not a whole number

OpenStudy (anonymous):

it would be a and c no c and d :/

OpenStudy (anonymous):

-1 and -2 are negative integers but -1-(-2)=1 is not a negative integer

OpenStudy (anonymous):

@JOHN__22 check my reasoning

OpenStudy (matlee):

it is a and c trust me im a major in math

OpenStudy (anonymous):

Hello @matlee read the book again and the question too

OpenStudy (anonymous):

0 and 1 are whole numbers but 0-1=-1 is not a whole number -1 and -2 are negative integers but -1-(-2)=1 is not a negative integer So the options A and B are incorrect

OpenStudy (anonymous):

http://openstudy.com/users/midhun.madhu1987

OpenStudy (anonymous):

Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. whole numbers B. negative integers C. integers D. rational numbers

OpenStudy (anonymous):

A. and C.

OpenStudy (anonymous):

C and D

OpenStudy (anonymous):

arenot you reading my reasoning??

OpenStudy (anonymous):

oh yeah xD

OpenStudy (anonymous):

i see your reasoning is correct

OpenStudy (anonymous):

and rational-rational is rational integer-integer is integer so the options C and D are correct

OpenStudy (anonymous):

@gatorboy91